# Efficient Simulation of Open Quantum Systems on NISQ Devices

### v1

@@ Abstract
Simulating open quantum systems, which interact with external environments, presents significant challenges on noisy intermediate-scale quantum (NISQ) devices due to limited qubit resources and noise. In this paper, we propose an efficient framework for simulating open quantum systems on NISQ hardware by leveraging a time-perturbative Kraus operator representation of the system’s dynamics. Our approach avoids the computationally expensive Trotterization method and exploits the Lindblad master equation to represent time evolution in a compact form, particularly for systems satisfying specific commutation relations. We demonstrate the efficiency of our method by simulating quantum channels, such as the continuous-time Pauli channel and damped harmonic oscillators, on NISQ devices, including IonQ Harmony and Quantinuum H1-1. Additionally, we introduce hardware-agnostic error mitigation techniques, including Pauli channel fitting and quantum depolarizing channel inversion, to enhance the fidelity of quantum simulations. Our results show strong agreement between the simulations on real quantum hardware and exact solutions, highlighting the potential of Kraus-based methods for scalable and accurate simulation of open quantum systems on NISQ devices. This framework opens pathways for simulating more complex systems under realistic conditions in the near term.
@@ I Introduction
In this work, we introduce more efficient methods for simulating open quantum systems, including novel circuit designs and error mitigation strategies. These approaches aim to push the boundaries of what is possible on NISQ devices and overcome some of the limitations of earlier methods. We focus on Kraus representation-based methods to simulate the dynamics of these systems. These methods exploit a representation of the system’s time evolution in terms of Kraus operators, which model the interaction between the system and its environment. When specific commutation relations between the system’s Hamiltonian and Lindblad operators are satisfied, this representation can be computed efficiently without Trotterization. The Kraus operators offer a more efficient and NISQ-friendly way of simulating these systems. We apply techniques to map the Kraus operators and time evolution operators to quantum circuits for implementation on actual NISQ hardware (IonQ Harmony and Quantinuum H1-1 device).
@@ II.2 Trotter-Based Methods
Solving (1) for arbitrary quantum systems as a closed-form function of time is known to be a difficult problem. As a result, one must commonly resort to numerical integration methods on quantum or classical computers. On quantum computers, the most popular of these methods are those that employ Trotter product formulas [15]. This is usually achieved by re-writing the right-hand side of (1) in the equivalent superoperator form, namely
@@
such that the evolution of each component $\mathcal{D}_{i}$ for a small time step $\delta t$ can be realized as an efficient sequence of quantum gates $U_{\mathcal{D}_{i}}\simeq\exp(\delta t\mathcal{D}_{i})$. The key idea of a Trotter product formula is that one can repeatedly apply an alternating sequence of the gates $U_{\mathcal{D}_{i}}$ over many discrete $\delta t$ time steps such that the approximation error vanishes in the limit as $\delta t\rightarrow 0$. For example, applying a second order Trotter product formula [16] to the decomposition (3) over an $n$-step time evolution period $t$ ($\delta t=t/n$), we obtain the approximation:
@@ II.3 Kraus Representation-Based Methods
For some systems, Trotterization can produce quantum circuits that require a large number of gates and qubits, which is undesirable for NISQ applications. However, for some special classes of systems, a closed-form solution to the Lindblad equation can be found that significantly reduces the complexity of Trotterization or even avoids Trotterization altogether. This closed form solution can always be written in the form of a time-dependent Kraus operator representation of the solution to (1), given by:
@@
then (1) can be solved in closed form by a time-perturbative Kraus series of the general form
@@
where $E$ is a real eigenvalue of $H$ and each $\xi_{n}$ is a non-negative real eigenvalue of $L_{n}^{\dagger}L_{n}$. Since the imaginary component of each eigenvalue $E_{\text{eff}}$ is non-positive, the effective Hamiltonian evolution operator $T(t)$ is a contraction, meaning $\left\lVert T(t)\rho T(t)^{\dagger}\right\rVert\leq\left\lVert\rho\right\rVert$ for any density matrix $\rho$ and time $t\geq 0$. The time-dependence of the Kraus operator terms in (7) is described by the positive function $f(t)$, which takes different forms, depending on the values of $\nu$ and $\lambda$ as defined in conditions (iii) and (iv). In the most general case, $f(x)$ takes the form
@@
where $\alpha=2\operatorname{Im}(\nu)/\hbar-\operatorname{Re}(\lambda)$ is a positive constant. When $\alpha=0$, we observe that the worst case time-dependence of the Kraus operators is such that $K_{m,\vec{k}}(t)\sim\sqrt{t^{m}/m!}$. This case corresponds to systems with unitary symmetries (e.g. Pauli channels), where the unitary symmetry operators are used as Lindblad operators with different choice of damping parameters $\gamma_{n}$. The $\alpha>0$ case corresponds to systems where $H$ and the Lindblad terms $L_{n}^{\dagger}L_{n}$ exhibit non-trivial commutation relations (e.g. bosonic systems with finite particle number). For these systems, the time dependence can be reduced to $K_{m,\vec{k}}(t)\sim\sqrt{1/m!}$, which means that the Kraus series converges much faster and exhibits much lower truncation error than the general case. For additional details regarding the derivation of the general form of the time-perturbative Kraus operators (7), we refer the reader to reference [21].
@@ II.3.1 Mapping to Quantum Circuits
Upon inspection of (6) and (7), we observe that if the effective time evolution operator $T(t)$ and the Lindblad operators $L_{n}$ can be efficiently realized as quantum circuits acting on a pure state |$\ket{\psi}$ (or a purified superposition $\sum_{i}\sqrt{p_{i}}\ket{\psi_{i}}$), then one can straightforwardly simulate the time-evolution of an arbitrary density matrix $\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ | on a quantum computer. This is achieved by first applying the sequence of Lindblad operators $\prod_{i}L_{\vec{k}_{i}}$ followed by $T(t)$, and then multiplying by the scalar term
@@
where $b_{n}=\left\lVert L_{n}\right\rVert$ is the normalization constant. This transformation has no effect on the dynamics of (1), and ensures that each $L_{n}$ can be encoded in a unitary of norm $1$. Once the $L_{n}$ operators are normalized, they must be expanded to unitaries that can be implemented with quantum gates. This is most commonly achieved through a unitary block-encoding scheme in which the transformed Lindblad operators are embedded in a block unitary $U_{L_{n}}$ that requires at least one clean ancilla qubit per application in a quantum circuit [22]. One such block form is given by the unitary Sz.-Nagy dilation [23] of $L_{n}$:
@@ II.3.2 Measurement and Readout of Circuits
For each operator (7) in the Kraus series expansion, the unitaries $U_{L_{n}}$ (implementing the Lindblad operators) and $T(t)$ (implementing the effective Hamiltonian evolution) can be combined to produce a circuit $U_{K_{m,\vec{k}}}(t)$ corresponding to each Kraus operator, as shown in Figure 1. When each circuit is run on a quantum device, both the system qubits and all ancillas are measured. When the ancillas all read |$\ket{0}$, the the resulting system qubits are measured in a desired basis and the estimated outcome probabilities for each circuit are added on a classical computer with scalar weights $|a_{m,\vec{k}}|^{2}$.
Figure 1: Diagram of the quantum circuit $U_{K_{m,\vec{k}}}(t)$ corresponding to a single Kraus operator in the time-perturbative Kraus series (6). This circuit applies a set of unitary dilations of the Lindblad operators ($U_{L_{\vec{k}}}$) indexed by the vector $\vec{k}$. Then, the effective Hamiltonian evolution $T(t)$ is applied. When all of the ancillas are measured to read |$\ket{0}$, the resulting measurement probabilities on the system qubits are re-scaled by the factor $|a_{m,\vec{k}}|^{2}$ and added on a classical computer.
@@ II.4.2 Bosonic Systems
where $M$ is a Hermitian matrix, $\Delta$ is a symmetric matrix, and the Lindblad operators are in terms of the transformed bosonic annihilation operator $L_{i}=\hat{a}_{i}$. This Hamiltonian (19) can realize the four wave-mixing/spontaneous parametric down conversion Hamiltonian [32] in a lossy cavity. This is particularly relevant in quantum imaging when trying to obtain strong squeezing in leaky cavities [33]. There has also been interest in using Lindbladians with dissipative couplings to engineer non-abelian states in photonic lattices [34]. The case where $\Delta_{jk}=0$ can be related to the dissipative evolution of various spin and fermionic Hamiltonians in the single-excitation subspace [35], as long as the Lindblad operators are particle/excitation number conserving.
@@ II.4.3 Lindblad Operators with Finite Group and Semigroup Structure
Although the finite sum of Kraus operators in (22) contains many individual terms, we observe that this series can be factored as the product of polynomial functions of the Lindblad operators. Specifically, in the finite abelian case with $\theta_{n}\neq 0$, we can re-scale the Lindblad operators via the transformation (13) with $b_{n}=\theta_{n}^{1/(2\ell_{n})}$. This allows us to write the time evolution of the system as $\rho(t)=S(t)\rho S(t)^{\dagger}$ where $S(t)$ is given by
@@
where $L_{n}^{\prime}=L_{n}/\theta^{1/(2\ell_{n})}$, $\gamma_{n}^{\prime}=\theta^{1/\ell_{n}}\gamma_{n}$, and $\Gamma^{\prime}=\sum_{n=1}^{N}\gamma_{n}^{\prime}$. If one can efficiently apply a controlled form of the gates $U_{L_{n}^{\prime}}$ that implement each re-scaled Lindblad operator $L_{n}^{\prime}$, then each factor in (25) can be applied using circuits of the form shown in Figure 2, provided that the distribution
@@ III.1 Continuous-Time Pauli Channel
where each $\sigma_{ij}\in\{I,X,Y,Z\}$ is a single-qubit Pauli operator and $N_{q}$ is the number of qubits. Since $H=0$ and the operators $L_{n}$ are unitary, we have $[H,L_{n}]=0$, which means that the system satisfies conditions (i)-(iv) as discussed in Section II.4.1. In addition, it can be shown that Pauli string Lindblad operators form a finite abelian group that satisfies properties (20) and (21) with $\ell_{n}=2$ for each $L_{n}$. This means that the system admits a finite Kraus representation of the form (22) with Kraus operators
@@
This operator can be efficiently simulated using a quantum circuit of the form shown in Figure 2. Since $\ell_{n}=2$, only a single ancilla qubit is needed to prepare the distribution |$\ket{\phi_{F_{n}}(t)}$ corresponding to each Lindblad operator and each evolution period $t$. This circuit requires one two-qubit gate for each Pauli component $\sigma_{ij}\neq I$ in the set of Lindblad operators. Alternatively, one can also simulate a Pauli channel using the non-factored Kraus operators in (28), making the observation that the time-dependent terms $e^{-\Gamma t/2}\prod_{n}\sqrt{F_{2,m_{n}}^{1}(\gamma_{n}t)}$ can be applied on a classical computer after measuring the initial state under the action of the powers $L_{n}^{m}$ of each Kraus operator. This requires performing partial (or in the worst case, a full) tomography on the initial state; however, the result of these measurements can be re-combined classically to find the continuous system trajectory for all values of $t>0$.
@@ III.2 Damped Schwinger Oscillator Model
Up to addition of scalar terms, this Hamiltonian can be written as a quadratic product of bosonic creation and annihilation operators of the form (17) with the matrix $M=\hbar\omega I$, where $I$ is the $2\times 2$ identity. Following the analysis in Section II.4.2, we can efficiently simulate quadratic Lindblad operators of the form (18), provided that the associated matrices $M$ and $V_{n}$ satisfy conditions (i)-(iv) in place of $H$ and $L_{n}$ respectively. For the $2$-mode case, a natural choice of $V_{n}$ is the set of Pauli operators ($X,Y,Z$), which generate $SU(2)$. In their bosonic quadratic form, these are proportional to the quantum angular momentum operators, as first proposed by Schwinger:
@@
Since $M=\hbar\omega I$ and the Pauli operators $V_{n}\in\{X,Y,Z\}$ meet the conditions for satisfying (i)-(iv), we can simulate the evolution of the system under a subset of the Lindblad operators (33)-(35) using a time-perturbative Kraus series of the form (6) with $f(t)=t$.
@@
In Figure 4 we plot the expectation value of the trajectory of the system in mass-independent coordinates using the operators $x_{0}=(\hat{a}_{1}^{\dagger}+\hat{a}_{1})/\sqrt{2},y_{0}=(\hat{a}_{2}^{\dagger}+\hat{a}_{2})/\sqrt{2}$ and their momentum analogues $p_{x0}=i(\hat{a}_{1}^{\dagger}-\hat{a}_{1})/\sqrt{2},p_{y0}=i(\hat{a}_{2}^{\dagger}-\hat{a}_{2})/\sqrt{2}$, where $\hat{a}_{1},\hat{a}_{2}$ are restricted to act on only the lowest modes |$\ket{0}$ and |$\ket{1}$. For both the position and momentum coordinates, we observe strong agreement between the exact solution and the estimated values (for both the noisy H1 device and a simulated ideal noiseless device).
@@
where $i,i^{\prime}$ are the coordinate (i.e. mode) indices and $n,n^{\prime}$ are the coordinate eigenstate indices. The analogous momentum density can be obtained through a Fourier transform of (37). In Figure 5 we plot the position and momentum density estimated on the Quantinuum H1-1 device for the first four points in time.

### v2

# Efficient Simulation of Open Quantum Systems on NISQ Trapped-Ion Hardware

@@ Abstract
Simulating open quantum systems, which interact with external environments, presents significant challenges on noisy intermediate-scale quantum (NISQ) devices due to limited qubit resources and noise. In this paper, we propose an efficient framework for simulating open quantum systems on NISQ hardware by leveraging a time-perturbative Kraus operator representation of the system’s dynamics. Our approach avoids the computationally expensive Trotterization method and exploits the Lindblad master equation to represent time evolution in a compact form, particularly for systems satisfying specific commutation relations. We demonstrate the efficiency of our method by simulating quantum channels, such as the continuous-time Pauli channel and damped harmonic oscillators, on NISQ trapped-ion hardware, including IonQ Harmony and Quantinuum H1-1. Additionally, we introduce hardware-agnostic error mitigation techniques, including Pauli channel fitting and quantum depolarizing channel inversion, to enhance the fidelity of quantum simulations. Our results show strong agreement between the simulations on real quantum hardware and exact solutions, highlighting the potential of Kraus-based methods for scalable and accurate simulation of open quantum systems on NISQ devices. This framework opens pathways for simulating more complex systems under realistic conditions in the near term.
@@ I Introduction
In this work, we introduce more efficient methods for simulating open quantum systems, including novel circuit designs and error mitigation strategies. These approaches aim to push the boundaries of what is possible on NISQ devices and overcome some of the limitations of earlier methods. We focus on Kraus representation-based methods to simulate the dynamics of these systems. These methods exploit a representation of the system’s time evolution in terms of Kraus operators, which model the interaction between the system and its environment. When specific commutation relations between the system’s Hamiltonian and Lindblad operators are satisfied, this representation can be computed efficiently without Trotterization. The Kraus operators offer a more efficient and NISQ-friendly way of simulating these systems. We apply techniques to map the Kraus operators and time evolution operators to quantum circuits for implementation on actual NISQ hardware (IonQ Harmony and Quantinuum H1-1 devices).
@@ II.2 Trotter-Based Methods
Solving Eq. (1) for arbitrary quantum systems as a closed-form function of time is known to be a difficult problem. As a result, one must commonly resort to numerical integration methods on quantum or classical computers. On quantum computers, the most popular of these methods are those that employ Trotter product formulas [15]. This is usually achieved by re-writing the right-hand side of Eq. (1) in the equivalent superoperator form, namely
@@
such that the evolution of each component $\mathcal{D}_{i}$ for a small time step $\delta t$ can be realized as an efficient sequence of quantum gates $U_{\mathcal{D}_{i}}\simeq\exp(\delta t\mathcal{D}_{i})$. The key idea of a Trotter product formula is that one can repeatedly apply an alternating sequence of the gates $U_{\mathcal{D}_{i}}$ over many discrete $\delta t$ time steps such that the approximation error vanishes in the limit as $\delta t\rightarrow 0$. For example, applying a second order Trotter product formula [16] to Eq. (3) over an $n$-step time evolution period $t$ ($\delta t=t/n$), we obtain the approximation:
@@ II.3 Kraus Representation-Based Methods
For some systems, Trotterization can produce quantum circuits that require a large number of gates and qubits, which is undesirable for NISQ applications. However, for some special classes of systems, a closed-form solution to the Lindblad equation can be found that significantly reduces the complexity of Trotterization or even avoids Trotterization altogether. This closed form solution can always be written in the form of a time-dependent Kraus operator representation of the solution to Eq. (1), given by:
@@
then Eq. (1) can be solved in closed form by a time-perturbative Kraus series of the general form
@@
where $E$ is a real eigenvalue of $H$ and each $\xi_{n}$ is a non-negative real eigenvalue of $L_{n}^{\dagger}L_{n}$. Since the imaginary component of each eigenvalue $E_{\text{eff}}$ is non-positive, the effective Hamiltonian evolution operator $T(t)$ is a contraction, meaning $\left\lVert T(t)\rho T(t)^{\dagger}\right\rVert\leq\left\lVert\rho\right\rVert$ for any density matrix $\rho$ and time $t\geq 0$. The time-dependence of the Kraus operator terms in Eq. (7) is described by the positive function $f(t)$, which takes different forms, depending on the values of $\nu$ and $\lambda$ as defined in conditions (iii) and (iv). In the most general case, $f(x)$ takes the form
@@
where $\alpha=2\operatorname{Im}(\nu)/\hbar-\operatorname{Re}(\lambda)$ is a positive constant. When $\alpha=0$, we observe that the worst case time-dependence of the Kraus operators is such that $K_{m,\vec{k}}(t)\sim\sqrt{t^{m}/m!}$. This case corresponds to systems with unitary symmetries (e.g. Pauli channels), where the unitary symmetry operators are used as Lindblad operators with different choice of damping parameters $\gamma_{n}$. The $\alpha>0$ case corresponds to systems where $H$ and the Lindblad terms $L_{n}^{\dagger}L_{n}$ exhibit non-trivial commutation relations (e.g. bosonic systems with finite particle number). For these systems, the time dependence can be reduced to $K_{m,\vec{k}}(t)\sim\sqrt{1/m!}$, which means that the Kraus series converges much faster and exhibits much lower truncation error than the general case. For additional details regarding the derivation of the general form of the time-perturbative Kraus operators in Eq. (7), we refer the reader to reference [21].
@@ II.3.1 Mapping to Quantum Circuits
Upon inspection of Eq. (6) and Eq. (7), we observe that if the effective time evolution operator $T(t)$ and the Lindblad operators $L_{n}$ can be efficiently realized as quantum circuits acting on a pure state |$\ket{\psi}$ (or a purified superposition $\sum_{i}\sqrt{p_{i}}\ket{\psi_{i}}$), then one can straightforwardly simulate the time-evolution of an arbitrary density matrix $\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ | on a quantum computer. This is achieved by first applying the sequence of Lindblad operators $\prod_{i}L_{\vec{k}_{i}}$ followed by $T(t)$, and then multiplying by the scalar term
@@
where $b_{n}=\left\lVert L_{n}\right\rVert$ is the normalization constant. This transformation has no effect on the dynamics of Eq. (1), and ensures that each $L_{n}$ can be encoded in a unitary of norm $1$. Once the $L_{n}$ operators are normalized, they must be expanded to unitaries that can be implemented with quantum gates. This is most commonly achieved through a unitary block-encoding scheme in which the transformed Lindblad operators are embedded in a block unitary $U_{L_{n}}$ that requires at least one clean ancilla qubit per application in a quantum circuit [22]. One such block form is given by the unitary Sz.-Nagy dilation [23] of $L_{n}$:
@@ II.3.2 Measurement and Readout of Circuits
For each operator Eq. (7) in the Kraus series expansion, the unitaries $U_{L_{n}}$ (implementing the Lindblad operators) and $T(t)$ (implementing the effective Hamiltonian evolution) can be combined to produce a circuit $U_{K_{m,\vec{k}}}(t)$ corresponding to each Kraus operator, as shown in Figure 1. When each circuit is run on a quantum device, both the system qubits and all ancillas are measured. When the ancillas all read |$\ket{0}$, the the resulting system qubits are measured in a desired basis and the estimated outcome probabilities for each circuit are added on a classical computer with scalar weights $|a_{m,\vec{k}}|^{2}$.
Figure 1: Diagram of the quantum circuit $U_{K_{m,\vec{k}}}(t)$ corresponding to a single Kraus operator in the time-perturbative Kraus series Eq. (6). This circuit applies a set of unitary dilations of the Lindblad operators ($U_{L_{\vec{k}}}$) indexed by the vector $\vec{k}$. Then, the effective Hamiltonian evolution $T(t)$ is applied. When all of the ancillas are measured to read |$\ket{0}$, the resulting measurement probabilities on the system qubits are re-scaled by the factor $|a_{m,\vec{k}}|^{2}$ and added on a classical computer.
@@ II.4.2 Bosonic Systems
where $M$ is a Hermitian matrix, $\Delta$ is a symmetric matrix, and the Lindblad operators are in terms of the transformed bosonic annihilation operator $L_{i}=\hat{a}_{i}$. This Hamiltonian in Eq. (19) can realize the four wave-mixing/spontaneous parametric down conversion Hamiltonian [32] in a lossy cavity. This is particularly relevant in quantum imaging when trying to obtain strong squeezing in leaky cavities [33]. There has also been interest in using Lindbladians with dissipative couplings to engineer non-abelian states in photonic lattices [34]. The case where $\Delta_{jk}=0$ can be related to the dissipative evolution of various spin and fermionic Hamiltonians in the single-excitation subspace [35], as long as the Lindblad operators are particle/excitation number conserving.
@@ II.4.3 Lindblad Operators with Finite Group and Semigroup Structure
Although the finite sum of Kraus operators in Eq. (22) contains many individual terms, we observe that this series can be factored as the product of polynomial functions of the Lindblad operators. Specifically, in the finite abelian case with $\theta_{n}\neq 0$, we can re-scale the Lindblad operators via the transformation in Eq. (13) with $b_{n}=\theta_{n}^{1/(2\ell_{n})}$. This allows us to write the time evolution of the system as $\rho(t)=S(t)\rho S(t)^{\dagger}$ where $S(t)$ is given by
@@
where $L_{n}^{\prime}=L_{n}/\theta^{1/(2\ell_{n})}$, $\gamma_{n}^{\prime}=\theta^{1/\ell_{n}}\gamma_{n}$, and $\Gamma^{\prime}=\sum_{n=1}^{N}\gamma_{n}^{\prime}$. If one can efficiently apply a controlled form of the gates $U_{L_{n}^{\prime}}$ that implement each re-scaled Lindblad operator $L_{n}^{\prime}$, then each factor in Eq. (25) can be applied using circuits of the form shown in Figure 2, provided that the distribution
@@ III.1 Continuous-Time Pauli Channel
where each $\sigma_{ij}\in\{I,X,Y,Z\}$ is a single-qubit Pauli operator and $N_{q}$ is the number of qubits. Since $H=0$ and the operators $L_{n}$ are unitary, we have $[H,L_{n}]=0$, which means that the system satisfies conditions (i)-(iv) as discussed in Section II.4.1. In addition, it can be shown that Pauli string Lindblad operators form a finite abelian group that satisfies Eq. (20) and Eq. (21) with $\ell_{n}=2$ for each $L_{n}$. This means that the system admits a finite Kraus representation of the form (22) with Kraus operators
@@
This operator can be efficiently simulated using a quantum circuit of the form shown in Figure 2. Since $\ell_{n}=2$, only a single ancilla qubit is needed to prepare the distribution |$\ket{\phi_{F_{n}}(t)}$ corresponding to each Lindblad operator and each evolution period $t$. This circuit requires one two-qubit gate for each Pauli component $\sigma_{ij}\neq I$ in the set of Lindblad operators. Alternatively, one can also simulate a Pauli channel using the non-factored Kraus operators in Eq. (28), making the observation that the time-dependent terms $e^{-\Gamma t/2}\prod_{n}\sqrt{F_{2,m_{n}}^{1}(\gamma_{n}t)}$ can be applied on a classical computer after measuring the initial state under the action of the powers $L_{n}^{m}$ of each Kraus operator. This requires performing partial (or in the worst case, a full) tomography on the initial state; however, the result of these measurements can be re-combined classically to find the continuous system trajectory for all values of $t>0$.
@@ III.2 Damped Schwinger Oscillator Model
Up to addition of scalar terms, this Hamiltonian can be written as a quadratic product of bosonic creation and annihilation operators of the form in Eq. (17) with the matrix $M=\hbar\omega I$, where $I$ is the $2\times 2$ identity. Following the analysis in Section II.4.2, we can efficiently simulate quadratic Lindblad operators of the form in Eq. (18), provided that the associated matrices $M$ and $V_{n}$ satisfy conditions (i)-(iv) in place of $H$ and $L_{n}$ respectively. For the $2$-mode case, a natural choice of $V_{n}$ is the set of Pauli operators ($X,Y,Z$), which generate $SU(2)$. In their bosonic quadratic form, these are proportional to the quantum angular momentum operators, as first proposed by Schwinger:
@@
Since $M=\hbar\omega I$ and the Pauli operators $V_{n}\in\{X,Y,Z\}$ meet the conditions for satisfying (i)-(iv), we can simulate the evolution of the system under a subset of the Lindblad operators Eqs. (33)-(35) using a time-perturbative Kraus series of the form in Eq. (6) with $f(t)=t$.
@@
In Figure 4 we plot the expectation value of the trajectory of the system in mass-independent coordinates using the operators $x_{0}=(\hat{a}_{1}^{\dagger}+\hat{a}_{1})/\sqrt{2},y_{0}=(\hat{a}_{2}^{\dagger}+\hat{a}_{2})/\sqrt{2}$ and their momentum analogues $p_{x0}=i(\hat{a}_{1}^{\dagger}-\hat{a}_{1})/\sqrt{2},p_{y0}=i(\hat{a}_{2}^{\dagger}-\hat{a}_{2})/\sqrt{2}$, where $\hat{a}_{1},\hat{a}_{2}$ are restricted to act on only the lowest modes |$\ket{0}$ and |$\ket{1}$. For both the position and momentum coordinates, we observe strong agreement between the exact solution and the estimated values (for both the noisy H1-1 device and a simulated ideal noiseless device).
@@
where $i,i^{\prime}$ are the coordinate (i.e. mode) indices and $n,n^{\prime}$ are the coordinate eigenstate indices. The analogous momentum density can be obtained through a Fourier transform of Eq. (37). In Figure 5 we plot the position and momentum density estimated on the Quantinuum H1-1 device for the first four points in time.